OpenStudy (anonymous):

check answer :) Rewrite the rational exponent as a radical by extending the properties of integer exponents. (more below)

OpenStudy (anonymous):

\[\frac{ 2^{\frac{7}{8}}}{ 2^{\frac{1}{4}} }\] I got \[\sqrt[8]{2^5}\]

Directrix (directrix):

I got that, too.

OpenStudy (anonymous):

because when it is set up like that you subtract the exponents to get 5/8, and the 5 goes under the radical while the 8 goes on the outside

OpenStudy (anonymous):

ok thanks its my first time doing radicals wanna make sure i start off good. i didn't start off great in functions

Directrix (directrix):

That "outside" is called the index. The index of your radical answer is 8.

Directrix (directrix):

You are off to a good start.

OpenStudy (anonymous):

oh ok, thanks!

OpenStudy (anonymous):

i really like radicals. not as hard as i thought when i did my pre-test

Directrix (directrix):

Keep practicing.